How do you quadratic formula q2 +6q-7=0

Answer by stanbon(75887)

How do you quadratic formula q2 +6q-7=0
----
a = 1 ; b = 6 ; c = -7
-----
Formula:
q = [-b +- sqrt(b^2-4ac)/(2a)
------
Your Problem:
q = [-6 +- sqrt(36-4*-7)]/2
----
q = [-6 +- sqrt(64)]/2
----
q = [-6 +- 8]/2
---
q = -14/2 or q = 2/2
---
q = -7 or q = 1
==================
Cheers,
Stan H.
==================